# Analyzing Uncertainty Quantification in Statistical and Deep Learning Models for Probabilistic Electricity Price Forecasting

### v2

@@ 1 Introduction
Despite various refined models, existing research often addresses data-related uncertainty, whereas uncertainty arising from model choice and calibration is often overlooked (see Section 2.1 for more details). Including this uncertainty, e.g., via ensemble or Bayesian approaches, considerably increases complexity and computational challenges, and it remains to be verified whether the additional effort pays off. So far, to the authors’ best knowledge, there has been no deeper analysis in electricity price forecasting with regard to this question. Hence, we propose to compare probabilistic forecasts and prediction intervals (PIs) generated by distributional deep learning and statistical models in combination with various uncertainty quantification methods. In addition, we discuss which evaluation measure is adequate here. For illustration and testing purposes, we use historical energy price data from the German market. These data not only show high volatility but also multiple structural breaks in the past and are therefore adequate for thoroughly testing methods for uncertainty quantification and comparison with forecasts from other models. In our forecasting study, we find that models from the related literature do not take model uncertainty into consideration. In this work, we demonstrate that accounting for both data and model uncertainty, particularly for deep learning models, yields significant benefits. As shown in Section 4, this approach leads to more reliable PIs as well as improved point forecasts. We also find that statistical models are highly competitive, even when compared to complex machine learning models.
The article is structured as follows: In Section 2, we discuss state-of-the-art approaches for (probabilistic/interval) electricity price forecasting from both machine learning and statistics, commenting on their approaches to quantifying uncertainty and motivating additional approaches to uncertainty quantification. In Section 3, the dataset for the case study is introduced. The forecasting study is presented in Section 4, where we discuss how models are calibrated and trained (including hyperparameter choices), and compare results using chosen metrics such as the prediction interval coverage probability (PICP) and the continuous ranked probability score (CRPS). Before we conclude the article in Section 6, we discuss our results in Section 5.
@@ 2 A Revision of Concepts for Forecasting and Uncertainty Quantification
There are various approaches for generating probabilistic and interval forecasts. First, there are parametric distribution-based approaches that rely on certain assumptions about the underlying probability distribution of the data. The PIs are constructed by estimating the parameters of the distribution and using the corresponding quantiles, as in distributional deep neural networks (DDNNs) [6]. Then, some methods directly learn specific aspects of a forecast, e.g., direct quantile learning methods such as quantile regression [7], which do not require distributional assumptions. Finally, there are model-agnostic approaches such as bootstrapping [8], historical simulation [9], or conformal prediction [9], which do not rely on any specific underlying distribution or regression model. These approaches are applied to already trained models.
@@ 2.2 Statistical Approaches
The standard approach to statistical interval forecasting is to assume an underlying distribution and obtain the interval by extracting the corresponding quantiles. The simplest way to model a distribution is the naive model. It uses the previous day’s price to forecast prices for Tuesday through Friday, and it uses prices from the same day of the previous week to predict prices for Monday, Saturday, and Sunday. To estimate price distributions, we apply historical simulation, relying on the widely used Gaussian distribution. This involves fitting distribution parameters to the forecast errors from a selected dataset. We refer to these models as Naive-HStrain and Naive-HSval, depending on whether the training or validation set was used.
With regard to the distribution’s volatility parameters, there are more complex approaches, such as the already mentioned generalized autoregressive conditional heteroskedasticity (GARCH) model, which proposes using an autoregressive model for volatility. Other distributions might be used as well. Authors such as Berrisch and Ziel [13] or Pappert and Arsova [14], for example, suggest using Copulas, which are a generalized distribution approach.
Here, we consider two competitive benchmark models for probabilistic forecasting, namely the LEAR-QRA model by [15], Sect. 4.2 (see also [16], Sect. 4.3.2), and the LEAR-GARCH model. Here, LEAR stands for LASSO-estimated autoregressive, and QRA for quantile regression averaging. In the following, both benchmark models are briefly described. We start by introducing the LEAR model for mean forecasting, which is the aspect both models share. Afterwards, the distributional modeling of the two approaches is described.
The LEAR model, originally introduced by [17], assumes an autoregressive model with external regressors for each product individually and then estimates the coefficients using a LASSO penalty [18]. For each product, the model may be written as
@@
The LEAR-GARCH model follows a different approach. Instead of relying on ensemble-based estimators for the probabilistic forecast, the conditional variance is predicted by a GARCH model. Assuming an underlying conditional distribution, the conditional mean (given by the LEAR prediction) and the conditional variance (given by the GARCH prediction) constitute a probabilistic forecast (in the application, we assume that the conditional distribution is the normal distribution).
While sharing a basic building block (the LEAR mean modeling), these two approaches are fundamentally different. The GARCH approach quantifies uncertainty explicitly by modeling the conditional variance, whereas the LEAR-QRA approach captures uncertainty implicitly using an ensemble of predictors trained on different horizons. In particular, short-term training windows can account for recent increases in variance, as discussed in [15].
@@ 2.3 Deep Learning-Based Concepts
As machine learning models are widely used today, we refrain from providing a detailed introduction to deep learning and machine learning models in general. If the reader is not familiar with such concepts, please see [19], for example.
Generally speaking, deep learning models learn complex patterns in data by processing it through several layers, thereby relying on the universal approximation theorem [20], which states that a feedforward deep neural network (FDNN) with enough neurons and a suitable activation function can approximate any continuous function defined on a closed and bounded subset of $\mathbb{R}^{n}$ with the desired precision. However, the theorem does not provide specific guidance on how to design the network, such as determining the optimal number of layers or neurons. Moreover, even if an FDNN is capable of representing a particular function, there is no guarantee that the training algorithm will successfully identify the appropriate parameters to achieve this approximation [21].
In modern NNs, the rectified linear unit (ReLU) activation function is the default choice for hidden layers, as recommended in [21]. However, the activation function used in the output layer depends on the specific problem being solved. For example, in regression problems such as price forecasting, a linear activation function is typically used. To summarize, designing an FDNN involves three key architectural choices: the number of layers, the number of neurons per layer, and the activation function used in each layer. In addition to FDNNs, several other NN architectures are used for electricity price forecasting, such as gated recurrent units and the temporal fusion transformer; see [22, 23] and references therein. These include recurrent neural networks (RNNs) and convolutional neural networks (CNNs), for example. RNNs are designed to process sequential data by using an internal memory called the hidden state, which allows them to retain information from previous steps [24]. The core operation in CNNs is convolution, which can be thought of as computing a weighted average at each input point by considering its surrounding $k$ neighboring values. These weights, known as kernels, help the model learn spatial or temporal patterns. A CNN typically consists of multiple kernels grouped within convolutional layers, which are trained using data [21].
After selecting an NN architecture, the next step is to train the model by learning its parameters. In the case of an FDNN, these parameters include the weights $W_{i}$ and biases $b_{i}$ for each layer $i=1\cdots L$. The goal of training is to adjust these parameters so that the network accurately captures the relationship between input features and output targets in a given dataset $\{x_{i},y_{i}\}_{i=1}^{N}$. This is achieved by minimizing a loss function. A commonly used loss function is the negative log-likelihood (NLL) loss, which is derived from the maximum likelihood estimation (MLE) method.
Here, we focus on DDNNs and ensemble models. A DDNN is a type of FDNN that predicts the parameters of a specific probability distribution. In this work, we use the normal distribution. The model is trained by minimizing the negative log-likelihood loss. To improve numerical stability during training, we follow the approach of [25] and predict the logarithm of the variance, $\log\sigma^{2}$, instead of the variance $\sigma^{2}$ directly.
@@ 2.3.1 Ensembles
The idea of ensemble methods is to combine different methods, i.e., their forecasts, to generate improved predictions. This, however, comes at a cost, as multiple models have to be calibrated and maintained. With the increasing application of machine learning tools in time series analysis, ensemble forecasts have also been introduced to this field [26]. Here, ensembles are often not combinations of different methods but rather several neural networks [27]. With regard to neural networks, ensemble models offer a way to evaluate uncertainty, as explained in [12] and [28]. This can be done by training different architectures, training on different subsets of the original dataset, or varying the initial parameter set. Here, we construct an ensemble model that consists of $N>0$ DDNNs and their outputs.
These are combined using a Gaussian mixture scheme. The respective PDF is constructed as a linear combination
$p(x;\theta)$ $=$ $\sum_{i=1}^{N}w_{i}\mathcal{N}(x;\mu_{i},\sigma_{i})$
using nonnegative weights $w_{1},\ldots,w_{N}$ that add up to one. The cumulative distribution function is defined accordingly, with $\theta=(\mu_{1},...,\mu_{N},\sigma_{1},...,\sigma_{N})$ and weights $w_{i},i=1,...,N$ [3]. To calculate the quantile $Q(p;\theta)=F^{-1}(p;\theta)$ for a given probability $p\in[0,1],$ we solve the following equation [29]:
@@
numerically. The initial guess for the solution is set to $x_{0}=\sum_{i=1}^{N}w_{i}F_{\mathcal{N}(\mu_{i},\sigma_{i})}^{-1}(p).$
2.3.2 Bayes - Monte Carlo Dropout
While ensemble models train a large number of NNs to describe model uncertainty, Bayesian logic suggests an alternative approach by interpreting individual weights as probabilistic variables. The Monte Carlo (MC) dropout model is a widely used example [30].
The Bayesian approach, MC Dropout, differs from the DDNN by including dropout layers after each hidden layer. These dropout layers remain active during both training and inference. As a result, the same input can produce different outputs in multiple forward passes through the network. During inference, we perform $N$ forward passes and aggregate the outputs using the Gaussian mixture scheme, equivalent to the ensemble method. Quantiles are computed by solving Equation 1, just as in the ensemble approach. The resulting NLL loss function using a Gaussian mixture distribution reads as follows:
$L_{NLL,GM}=-\sum_{i=1}^{N_{d,i}}\log\left(\sum_{j=1}^{N}\frac{1}{\sqrt{2\pi\hat{\sigma}_{i,j}^{2}}}\exp\left(-\frac{(y_{i}-\hat{\mu}_{i,j})^{2}}{2\hat{\sigma}_{i,j}^{2}}\right)\right)$
with predicted outputs $\hat{\mu}_{i,j}$ and $\hat{\sigma}_{i,j}$ of the NN for each forward pass. As with the DDNN model, we predict $\log\sigma^{2}$ instead of $\sigma^{2}$ directly.
@@ 2.4 Model-Agnostic Approaches
As already mentioned, these approaches do not require assumptions about a specific probability distribution or regression model. Instead, they can be applied on top of any forecasting model. Historical simulations are examples, as described in Section 2.2.
Another example is Conformal Prediction [9]. Given a trained model on historical data, one constructs PIs using nonconformity scores computed from the residuals of a distinct calibration set for each hour of the day. For each day $d$ and hour $h$, one computes the absolute residual as the nonconformity score:
@@ 3 Electricity Data
All data were retrieved from the SMARD platform (www.smard.de), operated by the German Federal Network Agency. For forecasting power prices, we consider not only the price itself but also load forecasts and renewable energy generation forecasts. Daily seasonality is captured using weekday dummies. The dataset covers the period from 1 October 2018 to 30 November 2024. In general, the data quality is very high. However, entire days are regularly missing in the load forecast series. The power prices we consider are hourly day-ahead prices with physical delivery, traded on the European Power Exchange (EPEX SPOT). The day-ahead auction is held daily, with price settlement taking place at noon for delivery on the following day.
@@ 4.2 Model Calibration and Training
All deep learning models share the same architecture, consisting of two hidden layers with 1,024 neurons each. The hidden layers use ReLU activation functions, while the output layer employs a linear activation. All deep learning models use the ”He” initialization for the hidden layers and the ”Glorot” initialization for the output layer. During training, early stopping is applied with a patience of 100 epochs, and the maximum number of training epochs is set to 2000. The models are optimized using the ADAM optimizer with a batch size of 32. To reduce overfitting, all models employ L2 regularization. Additionally, 10 independent runs are performed for each model, and the results are averaged. To ensure reproducibility, a fixed random seed is used for each training run. For the MCD10 and MCD30 models, only a single forward pass is performed during training, while validation involves 10 and 30 forward passes, respectively.
For the LEAR model, we follow [31] and train for different forecasting horizons, specifically 56, 84, 1092, and 1461 days. We average the results to produce the final point forecast. Additionally, we employ a rolling window scheme. The QRA and GARCH approaches use the mean forecasts obtained from the LEAR model based on the validation set. We fit 24 hourly GARCH models, each trained only on data from its corresponding hour.

### v3

@@ Abstract
Keywords: Bayesian methods, deep learning, electricity price forecasting, ensemble learning, interval forecasting, machine learning, Monte Carlo methods, power markets, probabilistic forecasting, uncertainty quantification
Preprint Submitted to IEEE Access
@@ 1 Introduction
Despite various refined models, existing research often addresses data-related uncertainty, whereas uncertainty arising from model choice and calibration is often overlooked (see Section 2.1 for more details). Including this uncertainty, e.g., via ensemble or Bayesian approaches, considerably increases complexity and computational challenges, and it remains to be verified whether the additional effort pays off. So far, to the authors’ best knowledge, there has been no deeper analysis in electricity price forecasting with regard to this question. Hence, we propose to compare probabilistic forecasts and prediction intervals (PIs) generated by distributional deep learning and statistical models in combination with various uncertainty quantification methods. In addition, we discuss which evaluation measure is adequate here. For illustration and testing purposes, we use historical energy price data from the German market. These data not only show high volatility but also multiple structural breaks in the past and are therefore adequate for thoroughly testing methods for uncertainty quantification and comparison with forecasts from other models. In this work, we demonstrate that accounting for both data and model uncertainty, particularly for deep learning models, yields significant benefits. As shown in Section 4, this approach leads to more reliable PIs as well as improved point forecasts. We also find that statistical models are highly competitive, even when compared to complex machine learning models.
The article is structured as follows: In Section 2, we discuss state-of-the-art approaches for (probabilistic/interval) electricity price forecasting from both machine learning and statistics, commenting on their approaches to quantifying uncertainty and motivating additional approaches to uncertainty quantification. In Section 3, the dataset for the case study is introduced. The forecasting study is presented in Section 4, where we discuss how models are calibrated and trained (including hyperparameter choices), and compare results using chosen metrics such as the prediction interval coverage probability (PICP) and the continuous ranked probability score (CRPS) (see [6]). Before we conclude the article in Section 6, we discuss our results in Section 5.
@@ 2 A Revision of Concepts for Forecasting and Uncertainty Quantification
There are various approaches for generating probabilistic and interval forecasts. First, there are parametric distribution-based approaches that rely on certain assumptions about the underlying probability distribution of the data. The PIs are constructed by estimating the parameters of the (conditional) distribution and using the corresponding quantiles, as in distributional deep neural networks (DDNNs) [7]. Then, some methods directly learn specific aspects of a forecast, e.g., direct quantile learning methods such as quantile regression [8], which do not require distributional assumptions. Finally, there are model-agnostic approaches such as bootstrapping [9], historical simulation [10], or conformal prediction [10], which do not rely on any specific underlying distribution or regression model. These approaches are applied to already trained models.
@@ 2.2 Statistical Approaches
The standard approach to statistical interval forecasting is to assume an underlying distribution and obtain the interval by extracting the corresponding quantiles. The simplest way to model a distribution is the naive model. The model naively uses a previous days price as mean forecast. To estimate price distributions, historical simulation may be applied, relying on the widely used Gaussian distribution. This involves fitting distribution parameters to the forecast errors from a selected dataset. We refer to these models as Naive-HStrain and Naive-HSval, depending on whether the training or validation set was used.
With regard to the distribution’s volatility parameters, there are more complex approaches, such as the generalized autoregressive conditional heteroskedasticity (GARCH) model (see [14]), which models the volatility parametrically, using an autoregressive model for it. Other parametric approaches such as copulas, see e.g. [15], might be used as well but are beyond the scope of this paper. The uncertainty may also be captured by ensemble methods or conformal prediction.
We consider three statistical models for probabilistic forecasting, namely the LEAR-QRA model by [2], Sect. 4.2 (see also [7], Sect. 4.3.2), the LEAR model with GARCH residuals, LEAR-GARCH, and the LEAR model with conformal prediction, LEAR-CP. LEAR stands for LASSO-estimated autoregressive, and QRA for quantile regression averaging. In the following, both benchmark models are briefly described. We start by introducing the LEAR model for mean forecasting, which is the aspect both models share. Afterwards, the distributional modeling of the two approaches is described.
The LEAR model, originally introduced by [16], assumes an autoregressive model with external regressors for each product individually and then estimates the coefficients using a LASSO penalty (see [17] for the LASSO penalty). For each product, the model may be written as
@@
The LEAR-GARCH model follows a different approach. Instead of relying on ensemble-based estimators for the probabilistic forecast, the conditional variance is predicted by a GARCH model. Assuming an underlying conditional distribution, the conditional mean (given by the LEAR prediction) and the conditional variance (given by the GARCH prediction) constitute a probabilistic forecast (in the application, we assume that the conditional distribution is the normal distribution). Last, in the LEAR-CP model, conformal prediction is applied to the mean-forecasts of the LEAR model.
While sharing a basic building block (the LEAR mean modeling), these approaches are fundamentally different. The GARCH approach quantifies uncertainty explicitly by modeling the conditional variance, whereas the LEAR-QRA approach captures uncertainty implicitly using an ensemble of predictors trained on different horizons. In particular, short-term training windows can account for recent increases in variance, as discussed in [2]. Then the conformal prediction approach offers a third, and different approach to uncertainty quantification, as is explained in the next section.
@@ 2.3 Deep Learning-Based Concepts
As machine learning models are widely used today, we refrain from providing a detailed introduction to deep learning and machine learning models in general. If the reader is not familiar with such concepts, we refer to [18].
Lago et al. [2] show that simple two-layer feed-forward neural networks are among the most effective deep learning models for point forecasting. Building on this, Marcjasz et al. [7] proposed the distributional deep neural network (DDNN), a probabilistic extension that estimates the parameters of the underlying price distribution rather than predicting electricity prices directly. The model is trained by minimizing the negative log-likelihood loss:
$L_{NLL}=-\sum_{i=1}^{N}\log p(y_{i}\mid x_{i};\theta),$
where $x_{i}$ and $y_{i}$ are the inputs and observed targets, and $p(y_{i}\mid x_{i};\theta)$ the probability assigned to $y_{i}$ by the distribution predicted by the network with parameters $\theta$. The output distribution captures data uncertainty in the predictions [11]. To additionally capture model uncertainty, various approaches can be used, including ensemble methods, Bayesian approaches such as Monte Carlo Dropout, and single deterministic methods such as Deep Evidential Regression [11].
@@ 2.3.1 Ensembles
The idea of ensemble methods is to combine different models, i.e., their forecasts, to generate improved predictions, which come at a cost, as multiple models must be calibrated and maintained.
However, ensembles also provide an intuitive way to measure model uncertainty by capturing the variation in predictions across the models. For NNs, it is essential to maximize the diversity of the network’s behavior to obtain reliable uncertainty estimates (while maintaining good predictive performance). This can be done by training different architectures, training on different subsets of the original dataset, or varying the initial parameter set [19]. Lakshminarayanan et al. [20] found that using different random initializations of the NN weights and random shuffling suffices to create network diversity. To get the predictive distribution, one has to aggregate the ensemble models into a weighted mixture model:
$p(y\mid x;\theta)$ $=$ $\sum_{j=1}^{M}w_{j}\,p(y\mid x;\theta_{j})$
using nonnegative weights $w_{1},\ldots,w_{M}$ that add up to one. To calculate the quantile $Q(p;\theta)=F^{-1}(p;\theta)$ for a given probability $p\in[0,1],$ we solve the following equation:
@@
numerically. The initial guess for the solution we set to $x_{0}=\sum_{i=1}^{M}w_{j}F_{j}^{-1}(p).$
2.3.2 Monte Carlo Dropout
While ensemble models train several NNs to capture model uncertainty, Bayesian logic suggests an alternative approach by interpreting individual weights as probabilistic variables. The Monte Carlo (MC) dropout model is a widely used example [21]. It differs from the DDNN by including dropout layers after each hidden layer. These dropout layers remain active during both training and inference. As a result, the same input can produce different outputs in multiple forward passes through the network. During inference, we perform $N$ forward passes and aggregate the results using a uniformly weighted mixture model. Quantiles are then obtained by solving Equation 1, equivalent to the ensemble approach.
2.3.3 Deep Evidential Regression
Deep evidential regression (DER) is a single deterministic method, introduced by Amini et al. [22], where one evidential deep neural network (EvDNN) predicts both the forecast and its uncertainty. In contrast to a Bayesian method, priors are placed directly over the parameters of the distribution instead of the network weights. If a Gaussian distribution is assumed, a Gaussian prior is placed on the mean and an inverse gamma prior on the variance:
$\begin{split}y&\sim\mathcal{N}(\mu,\sigma^{2}),\\ \text{with }\mu&\sim\mathcal{N}(\gamma,\sigma^{2}\nu^{-1}),\quad\sigma^{2}\sim\Gamma^{-1}(\alpha,\beta),\end{split}$
where $\gamma\in\mathbb{R}$, $\nu>0$, $\alpha>1$, and $\beta>0$. This leads to a normal-inverse gamma distribution for the joint prior distribution, assuming that the parameters factorize as $p(\mu,\sigma^{2})=p(\mu)(\sigma^{2})$:
$\begin{split}&p(\theta|m)\\ &=\dfrac{\beta^{\alpha}\sqrt{\nu}}{\Gamma(\alpha)\sqrt{2\pi\sigma^{2}}}\left(\dfrac{1}{\sigma^{2}}\right)^{\alpha+1}\exp\left(-\dfrac{2\beta+\nu(\gamma-\mu)^{2}}{2\sigma^{2}}\right)\\ &=NIG(\mu,\sigma^{2};\gamma,\nu,\alpha,\beta)\end{split},$
where we write $\theta=(\mu,\sigma^{2})$ and $m=\gamma,\nu,\alpha,\beta$. Integrating over $\theta$ yields the predictive distribution for $y$:
$\begin{split}p(y|m)&=\int p(y|\theta)p(\theta|m)\,d\theta\\ &=St\left(y;\gamma,\dfrac{\beta(1+\nu)}{\nu\alpha};2\alpha\right).\end{split}$
This predictive distribution follows a Student’s t-distribution with location $\gamma$, scale $\tfrac{\beta(1+\nu)}{\nu\alpha}$, and $2\alpha$ degrees of freedom. The network is then trained by minimizing the NLL loss
$\begin{split}L_{DER,NLL}&=\sum_{i=1}^{N}St\left(y_{i};\hat{\gamma}_{i},\dfrac{\hat{\beta}_{i}(1+\hat{\nu}_{i})}{\hat{\nu}_{i}\hat{\alpha}_{i}};2\hat{\alpha}_{i}\right)\\ &=\sum_{i=1}^{N}\dfrac{1}{2}\log\left(\dfrac{\pi}{\hat{\nu}_{i}}\right)-\hat{\alpha}_{i}\log(\hat{\Omega}_{i})\\ &+\left(\hat{\alpha}_{i}+\dfrac{1}{2}\right)\log\left((y-\hat{\gamma}_{i})^{2}\hat{\nu}_{i}+\hat{\Omega}_{i}\right)\\ &+\log\left(\dfrac{\Gamma(\hat{\alpha}_{i})}{\Gamma\left(\hat{\alpha}_{i}+\frac{1}{2}\right)}\right),\end{split}$
where $\hat{\Omega}_{i}=2\hat{\beta}_{i}(1+\hat{\nu}_{i})$, $\Gamma(\cdot)$ denotes the Gamma function, and $\hat{\gamma}_{i}$, $\hat{\nu}_{i}$, $\hat{\alpha}_{i}$, and $\hat{\beta}_{i}$ are the predicted parameters for each training sample $(x_{i},y_{i})$. In addition to the NLL loss, an evidential regularizer is introduced to penalize predictions that are inaccurate while exhibiting low uncertainty. This term is defined as
$L_{DER,R}=\sum_{i=1}^{N}|y_{i}-\hat{\gamma}_{i}|\cdot(2\hat{\nu}_{i}+\hat{\alpha}_{i}).$
The overall loss function for the DER method is the sum of the NLL loss and the regularizer, weighted by a hyperparameter $\lambda_{DER}>0$:
$L_{DER}=L_{DER,NLL}+\lambda_{DER}\,L_{DER,R}.$
@@ 2.4 Model-Agnostic Approaches
Model-agnostic approaches do not require assumptions about a specific probability distribution or regression model. Instead, they can be applied on top of any forecasting model. Historical simulations are examples, as described in Section 2.2.
Another example is Conformal Prediction [10]. Given a trained model on historical data, PIs are constructed using nonconformity scores, computed from the residuals of a distinct calibration set for each hour of the day. For each day $d$ and hour $h$, the absolute residual as the nonconformity score is computed as,
@@ 3 Electricity Data
All data were retrieved from the SMARD platform (smard.de), operated by the German Federal Network Agency. For forecasting power prices, we consider not only the price itself but also load forecasts and renewable energy generation forecasts. Daily seasonality is captured using weekday dummies. The dataset covers the period from 1 October 2018 to 30 November 2024. In general, the data quality is very high. However, entire days are regularly missing in the load forecast series. The power prices we consider are hourly day-ahead prices with physical delivery, traded on the European Power Exchange (EPEX SPOT). The day-ahead auction is held daily, with price settlement taking place at noon for delivery on the following day [7].
@@ 4.2 Model Calibration and Training
All deep learning models share the same architecture, consisting of two hidden layers with 1,024 neurons each. The hidden layers use ReLU activation functions, while the output layer employs a linear activation. All deep learning models use the ”He” initialization for the hidden layers and the ”Glorot” initialization for the output layer. During training, early stopping is applied with a patience of 100 epochs, and the maximum number of training epochs is set to 2000. The models are optimized using the ADAM optimizer with a batch size of 32. To reduce overfitting, all models employ L2 regularization. Additionally, 10 independent runs are performed for each model, and the results are averaged. To ensure reproducibility, a fixed random seed is used for each training run. For the MCD10 and MCD30 models, only a single forward pass is performed during training, while validation involves 10 and 30 forward passes, respectively. To improve numerical stability during training, we follow the approach of [23] and predict the logarithm of the variance, $\log\sigma^{2}$, instead of predicting the variance $\sigma^{2}$ directly. For the EvDNN, we similarly predict $\log\nu$, $\log(\alpha-1)$, and $\log\beta$ instead of $\nu$, $\alpha$, and $\beta$, and for the DDNN we predict $\log\sigma^{2}$ in place of $\sigma^{2}$.
For the LEAR model, we follow [3] and train for different forecasting horizons, specifically 56, 84, 1092, and 1461 days. We average the results to produce the final point forecast. Additionally, we employ a rolling window scheme. The QRA and GARCH approaches use the mean forecasts obtained from the LEAR model based on the validation set. We fit 24 hourly GARCH models, each trained only on data from its corresponding hour. For the naive model, we use the previous day’s price as the mean forecast for days during Tuesday and Friday, and prices from the same day of the previous week to predict prices for Monday, Saturday, and Sunday.